Streamsong Credit Bank is offering 6.4 percent compounded daily on its savings accounts. Assume that you deposit $6,100 today. Use 365 days in a year.

a. How much will you have in the account in 3 years?

b. How much will you have in the account in 6 years?

c. How much will you have in the account in 12 years?
Answer #1

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.A=6100*(1+0.064/365)^(365*3)

=6100*1.21165012

=$7391.07(Approx).

b.A=6100*(1+0.064/365)^(365*6)

=6100*1.46809602

=$8955.39(Approx).

c.A=6100*(1+0.064/365)^(365*12)

=6100*2.15530593

=$13147.37(Approx).